If the density of water is 1,00 g/cm° and atmospheric pressure is 1,013 x 105 Pa, the depth in water (in m) where the absolute pressure is four times atmospheric

pressure, is equal to:


1
Expert's answer
2022-06-02T10:25:56-0400
"p=p_a+g\rho h"

"4p_a=p_a+g\rho h"

"h=\frac{3p_a}{g\rho}"

"h=\frac{3*1.013*10^5}{9.8*1000}=31\:\rm m"